what is the slope for 11y + 5x = 14?

Found 2 solutions by rfer, Alan3354:
Answer by rfer(16322)

y=-5/11x+14/11
slope=-5/11

Answer by Alan3354(69443)

what is the slope for 11y + 5x = 14?
--------------
Solve for y, then the slope, m = the coefficient of x
----
It's the "slope of" not the "slope for"